Express (5.07x10^5) g in SI base units using scientific notation.
1
Expert's answer
2020-12-09T06:33:38-0500

S.I unit of mass is kg


To convert gram to kilogram, you divide gram by 1000

Therefore, "5.07\u00d710^5"g = "\dfrac{5.07×10^5}{1000} kg" = 5.07 × 10² kg


This means that 5.07×10⁵ g expressed in SI unit is 5.07×10² kg.